# CoStruction: Conjoint radiance field optimization for urban scene reconStruction with limited image overlap

### v1

@@ 3.2 Guided ray sampling
$[t_{n},t_{f}](r)=\begin{cases}\left[0,D_{\mathcal{E}}(r)+\delta\right],&\text{if }\mathcal{M}_{c}(r)=0\\ \left[0,\infty\right[,&\text{otherwise }\end{cases}$ (6)
where $\delta$ is a hyperparameter (analogous to the shell factor in [38]) periodically updated.
With this formulation, when rays intersect the mesh with a high confidence value, points are sampled near the mesh surface. The rationale behind this approach is that if the color on the mesh surface closely matches the ground truth (GT) color, it is likely that the mesh has converged to the true surface of the scene. In such a case, we can rely on the SDF representation to guide the volumetric representation.
GRS for SDF representation.
In contrast to the GRS of $\mathcal{F}_{\Theta}^{v}$, we leverage the uncertainty mask $\mathcal{M}_{d}$ to guide the SDF model. Similarly, we modify the near and far plane $[t_{n},t_{f}]$ of ray $r$ according to $\mathcal{M}_{d}$ :
$[t_{n},t_{f}](r)=\begin{cases}[D_{v}(r)-\delta,D_{v}(r)+\delta],&\text{if }\mathcal{M}_{d}(r)=0\\ [D_{\mathcal{E}}(r)-\delta,D_{\mathcal{E}}(r)+\delta],&\text{otherwise}\end{cases}$ (7)
@@
$\mathcal{M}_{d}(r)$ indicates regions where the two representations are consistent with each other, suggesting minimal geometric ambiguity. In these regions, we can trust the estimated geometry from both representations and concentrate sampling near the surface. Conversely, it is more subtle to determine which representation to trust in regions where predictions from the two representations are inconsistent. This inconsistency may occur if the volumetric representation fails on large texture-less planes or if the SDF representation fails to reconstruct fine structural details. In such a case, we propose a simple approach: prioritize the volumetric representation and refine any introduced noise through geometric regularization (see Sec. 3.3).
Adaptive thresholding
The ray classification step of the guided ray sampling strategy involves choosing thresholds $\tau_{d}$ and $\tau_{c}$ to distinguish between certain and uncertain rays. Selecting such threshold empirically can be complicated as for different scene and reconstruction, this value may change. We propose a simple yet effective adaptive threshold selection method to overcome this limitation. Starting from an initial threshold value, we compute on a regular interval during the reconstruction process the ratio between the number of certain and uncertain rays. We define a scheduling policy that is design to control this ratio along the whole training and we modify the threshold accordingly to meet the ratio requirement. For instance, if the at a certain step of the reconstruction process, the scheduling policy imposes a ratio of 20% of certain rays and 80% of uncertain rays and the current ratio is 30% of certain rays and 70% of uncertain rays, we will lower the threshold value to reduce the proportion of certain rays.
@@ 3.3 Optimization details
While our GRS modify the sampling boundaries of each ray according to uncertainty masks, we also use a density estimator to further refine the point sampling step. We use a proposal network similar to [9] that is self-supervised by $\mathcal{F}_{\Theta}^{v}$ using the proposal loss $\mathcal{L}_{p}$ introduced in [1].
@@
For both representation, $\mathcal{F}_{\Theta}^{v}$ AND $\mathcal{F}_{\Theta}^{f}$, we use the standard $L_{1}$ loss to minimize the pixel-wise color difference between the rendered image $\hat{C}$, and the ground truth image $C$, as well as the DSSIM [31] loss on color patches, denoted together as $\mathcal{L}_{\text{rgb}}$. We also apply distortion regularization $\mathcal{L}_{\text{d}}$ introduced in MipNeRF [1] to avoid floaters. Similar to StreetSurf [12], we model the sky color with an auxiliary MLP conditioned on the ray direction and use a sky loss $\mathcal{L}_{\text{sky}}$ to constrain the opacity of sky pixels to zero. The segmentation mask is obtained with an off-the-shelf semantic segmentation network [6]. Following SCILLA[9], we supervise the normal associated with the closest sample to the surface with pseudo-ground truth obtained from off-the-shelf network [10] with loss $\mathcal{L}_{\hat{N}}$.
To facilitate more complete fine structure learning, we add an additional semantic head to $\mathcal{F}_{\Theta}^{v}$ and apply an additional cross-entropy loss semantic supervision $\mathcal{L}_{\text{rgb}}$. To constrain our SDF model to respect the eikonal equation, we also apply a masked Eikonal loss $\mathcal{L}_{\text{eik}}$ described in the following paragraph to the SDF Field.
@@
$\mathcal{L}_{\text{f}}=\mathcal{L}_{\text{rgb}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{\text{sky}}\mathcal{L}_{\text{sky}}+\lambda_{\hat{N}}\mathcal{L}_{\hat{N}}+\lambda_{\text{eik}}\mathcal{L}_{\text{eik}}^{m}.$ (9)
where $\lambda_{d}$, $\lambda_{\text{sky}}$, $\lambda_{\text{normal}}$, $\lambda_{d}$, $\lambda_{\text{normal}}$, $\lambda_{\text{eik}}$ are scaling factors.
Adaptive masked Eikonal constrain
To facilitate fine structure learning at early stage, we take inspiration from RaNeuS [36] method and relax the eikonal $\mathcal{L}_{\text{eik}}$, together with the normal loss $\mathcal{L}_{\text{normal}}$ with the uncertainty mask $\mathcal{M}_{d}$:
$\mathcal{L}_{\text{eik}}^{m}=\frac{1}{mn}\sum_{i=1}^{m}(1-\mathcal{M}_{d}(r_{i}))\cdot\big{(}\sum_{j=1}^{n}\left(\|\nabla f(x_{(i,j)}\|_{2}-1\right)^{2}\big{)}.$ (10)
Here, $\mathbf{n}_{ij}=\nabla f(x_{(i,j)}$ is the gradient of the sign distance function $f$ of at sample $x_{(i,j)}$. The masked provides two key benefits. Firstly, at each training step, the Eikonal loss is applied only to regions where the SDF field is confident, thereby relaxing constraints on areas with missing structures to facilitate learning. Conversely, to later maintain the properties of a signed distance function, the loss will be applied more broadly once complex structures are captured by the SDF field. This adaptive approach is enabled by the mask calculation, which automatically detects changes in the field’s confidence as training progresses. We also mask the normal loss $\mathcal{L}_{\text{normal}}$ in the same manner to avoid over-regularization at early stage.
SDF refinement.
Following Wang et al. [35], we refine the SDF field at the end of the training with this two modifications:
  • we enhance the mononormal supervision along with stronger distortion regularization.
  • we replace the density estimation provided by the proposal network by computing PDF weight along the ray directly with $\mathcal{F}_{\Theta}^{f}$.
@@ 4 Experiments
We use hash encoding to encode the positions [23], and spherical harmonics to encode the viewing directions. We use 2 layers with 64 hidden units for the MLPs $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$. We traine our model on a single high-tier GPU using Adam optimizer with a cosine learning rate decay from $10^{-2}$ to $10^{-4}$. The average training time is around 40 mins per scenes. We use Marching Cubes [19] to generate the final mesh that represents the scene. Further implementation details can be found in the supplementary materials.
@@
[…]
@@
[…]
@@
We evaluate our method on four public driving datasets: KITTI-360 [18], nuScenes [2], Waymo Open Dataset [29] and Pandaset [39].We conduct our experiments on four scenes from each dataset with large differences. We conduct our experiments on static sequences and mask dynamic vehicles for Pandaset scenes.
@@
We compare our proposed solution to current state-of-the-art (SoTA) surface reconstruction methods, including the SDF-based approaches StreetSurf [12] and SCILLA [9], as well as the Gaussian splatting-based method GoF [42]. StreetSurf [12] models close, far, and sky regions using hierarchical space partitioning, with 3D and 4D hash grids and occupancy grids for efficient ray sampling. SCILLA [9] initializes the SDF field with volumetric density predictions, achieving SoTA results in autonomous driving scenario. GoF [42], a Gaussian splatting method, achieves top performance in object-centric scenes with high image overlap, using 2D Gaussian splatting regularization losses [14], and propose a better mesh extraction solution tailored for 3DGS. We train two variants of GoF: GoF-sparse, initialized from a COLMAP sparse point cloud, and GoF-dense, initialized from a COLMAP dense point cloud.
@@
Similar to SCILLA [9], we evaluate the quality of the reconstructed meshes with two metrics:
@@
Precision (Prec.): the percentage of LiDAR points with a distance to the mesh below $0.15$m.
StreetSurf [12] SCILLA [9] GoF [42]-dense CoStruction (ours)
@@
Figure 3: Qualitative experiments results on four popular autonomous driving datasets. Complex scene geometries reconstructed by the mesh are color highlighted. We compare our mesh to the ones from StreetSurf [12], SCILLA [9] and GoF [42]
4.1 Results
Quantitative analysis.
We report quantitative results across four datasets in Tab. 1. Our method consistently outperforms or matches state-of-the-art approaches, achieving the top metrics in most scenes on KITTI-360 and nuScenes and delivering competitive results on Pandaset and Waymo.
[…]
Our performance on Pandaset and Waymo datasets is comparatively lower. This is likely due to the significant occlusions caused by vegetation and trees, which are abundant in this dataset. Given our guided sampling approach, our method is sensitive to heavily occluded objects, leading to reduced accuracy in such scenes.
[…]
@@ 4.1 Results
To complete our evaluation, we further present qualitative results in Fig. 3 across different dataset with complex scenes geometry, including a mixture of large road surfaces and fine structures such as poles, traffic lights, and tree trunks..
Both SCILLA [9] and StreetSurf [12] are able to reconstruct smooth complete surface but result in incomplete fine structure reconstruction. GoF [42] produce hight quality on fine structures but fails large road plane reconstruction (Nuscenes 0064). Our results demonstrate that our method outperforms others in accurately reconstructing both surfaces and fine structures, under conditions of limited image overlap.
@@ 4.2 Ablation study
To have a clear understanding the contribution of each key component in our method, we conducted an ablation study on guided sampling directed by $\mathcal{M}_{c}$, masked regularization using $\mathcal{M}_{c}$ and proposal guidance during early stage. Quantitative and qualitative results from KITTI-360 dateset are shown in Tab. 2 and Fig. 4.
First, we observe that guided sampling significantly improves overall geometry, enhancing both surface accuracy and fine structural details. This is shown by metrics for the entire scene and specifically for thin structures, such as poles, as well as in the visual quality of fine details shown in Fig. 4. These results indicate that guided sampling, enabled by cross-representation uncertainty masks $\mathcal{M}_{d}$ and $\mathcal{M}_{c}$ effectively filter ambiguity under limited image overlap condition, leading to more accurate reconstructions.
@@
Lastly, proposal supervision in the initial training stage aids in completing thin structures by providing a stable foundation for complex details that are challenging to reconstruct.
Together, these components enable our method to achieve high-fidelity reconstructions in complex urban scenes with limited image overlap.
(a) w/o guided sampling
(b) w/o masked reg.
(c) w/o proposal sup.
(d) CoStruction (ours)
Figure 4: Ablations study: we deactivate some CoStruction ’s key components. (a) without guided ray sampling directed by $\mathcal{M}_{d,c}$ (b) without masking $\mathcal{L}_{\text{eik}}$ and $\mathcal{L}_{\text{normal}}$ with $\mathcal{M}_{c}$ (c) without proposal guidance for SDF initialization .
@@ 5 Conclusion
In this work, we presented CoStruction, a hybrid solution with conjoint radiance field optimization designed to reconstruct accurate geometry for complex urban scenes with limited view overlap. We introduced Guided Ray Sampling, directed by cross-representation uncertainty, to filter geometric ambiguities arising from partial observations and achieve precise geometry reconstruction in outdoor environments. Extensive quantitative and qualitative evaluations show that CoStruction outperforms state-of-the-art SDF and Gaussian splatting methods, delivering more accurate reconstructions on both large planar regions and fine structures under autonomous driving setting.

### v2

@@ 3.2 Guided Ray Sampling (GRS)
$t_{n},t_{f}](\mathbf{r})=\begin{cases}\left[0,D_{\mathcal{E}}(\mathbf{r})+\delta\right]&\text{if }\mathcal{\mu}_{c}<\tau_{c},\\ \left[0,\infty\right[&\text{otherwise, }\end{cases}$ (6)
where $\delta$ is a hyperparameter (analogous to the shell factor in [38]) periodically updated. This helps avoid noisy planar predictions that may arise from the under-constrained optimization of the volumetric representation, ensuring more accurate surface reconstruction. In practice, $\tau_{c}$ is set as a constant for each dataset (see Section. 2 in supplementary material).
SDF-GRS.
After rendering $\mathcal{F}_{\Theta}^{v}$ with guided sampling, the predicted depth $\hat{D}_{v}(\mathbf{r})$ becomes available, allowing us to estimate the geometric uncertainty $\mathcal{\mu}_{d}$ to direct the sampling of $\mathcal{F}_{\Theta}^{f}$. Since high $\mathcal{\mu}_{d}$ values indicates regions where the predictions of the two representations are inconsistent, it is crucial to decide which representation to trust. Here, we propose trusting the volumetric representation, as it typically learns faster and captures complex structures more effectively and expect that noisy predictions from $\mathcal{F}_{\Theta}^{v}$ will be handled by the Volumetric-GRS.
Similar to Volumetric-GRS, we adjust the sampling bounds based on the geometric uncertainty $\mathcal{\mu}_{d}(\mathbf{r})$. If $\mathcal{\mu}_{d}(\mathbf{r})$ exceeds a threshold $\tau_{d}$, we focus the sampling on the estimated depth from the volumetric field $\hat{D}_{v}(\mathbf{r})$. Otherwise, sampling is concentrated around the depth of the mesh $D_{\mathcal{E}}(\mathbf{r})$, ensuring accurate surface reconstruction in both cases:
$t_{n},t_{f}](\mathbf{r})=\begin{cases}[D_{\mathcal{E}}(\mathbf{r})-\delta,D_{\mathcal{E}}(\mathbf{r})+\delta]&\text{if }\mathcal{\mu}_{d}<\tau_{d},\\ [D_{v}(\mathbf{r})-\delta,D_{v}(\mathbf{r})+\delta]&\text{otherwise,}\end{cases}$ (7)
@@
Adaptive thresholding.
The SDF-GRS strategy employs a threshold $\tau_{d}$ to classify rays as “certain” or “uncertain,” but selecting an appropriate value a priori can be difficult across varying scenes and reconstruction tasks. To overcome this, we propose a lightweight, data‐driven algorithm that adaptively adjusts $\tau_{d}$ based on the observed ratio of certain to uncertain rays (see Algorithm 1). Empirically, this adaptive scheme generalizes robustly across diverse urban environments (see supplementary material for details).
Input: $\tau$, $\{\mu_{d}(r_{i})\}_{i=1}^{N}$: depth uncertainties of $N$ rays
Data: $\gamma_{\uparrow},\gamma_{\downarrow}$: growth/decay factors; $\rho_{\text{high}},\rho_{\text{low}}$: ratio thresholds
Output: Updated $\tau$
$u\leftarrow\sum_{i}[\mu_{d}(r_{i})>\tau]$; $c\leftarrow(N-u)$; $\rho\leftarrow(u/c)$;
if $\rho>\rho_{\text{high}}$ then
  $\tau\leftarrow\tau\times\gamma_{\uparrow}$;
else if $\rho<\rho_{\text{low}}$ then
  $\tau\leftarrow\tau\times\gamma_{\downarrow}$;
return $\tau$
Algorithm 1 Adaptive Uncertainty Threshold $\tau$
3.3 Adaptive Relaxation on Geometric Constrain
Another critical factor in capturing fine structures within the SDF field is avoiding over-regularization while the geometry remains under-optimized [36, 35]. Whereas previous urban reconstruction methods enforce the Eikonal constraint across the entire scene [9, 12], we adaptively relax it in regions where the SDF prediction is uncertain. Additionally, we relax normal supervision in the same way based on pseudo ground truth normals $\bar{N}(r)$ predicted by an off-the-shelf network [10], circumventing unreliable supervision. Below, we define our two uncertainty-aware geometric regularization term:
$\mathcal{L}_{N}^{u}=\mathbb{I}_{\mu_{c}}\cdot\Biggl(\Bigl\|\frac{\nabla f(x_{N})}{\|\nabla f(x_{N})\|_{2}}-\bar{N}(r)\Bigr\|\;+\;\\ \Bigl\|1-\left(\frac{\nabla f(x_{N})}{\|\nabla f(x_{N})\|_{2}}\right)^{\top}\bar{N}(r)\Bigr\|\Biggr),$ (8)
$\mathcal{L}_{\text{eik}}^{u}=\mathbb{I}_{\mu_{c}}\cdot\left(\left\|\nabla f(x_{(i,j)})\right\|_{2}-1\right)^{2},$ (9)
where $x_{N}$ denotes the closest sample to the surface, as described in [9]. We define the indicator function $\mathbb{I}_{\mu_{c}}$ as follows, which can be regarded as a trimmed robust kernel that down-weights uncertain regions during supervised loss computation, in line with RobustNeRF [28]:
$\mathbb{I}_{\mu_{c}}\;=\;\begin{cases}1,&\text{if }\mu_{c}(\mathbf{r})>\tau_{c},\\[6.0pt] 0,&\text{otherwise}.\end{cases}$
@@ 3.4 Optimization details
While the GRS mechanism adjusts the sampling boundaries of each ray based on uncertainty, we further refine the sampling process using a density estimator. Specifically, we employ a proposal network – similar to ViiNeuS [9] – that is self-supervised by the volumetric field $\mathcal{F}_{\Theta}^{v}$ using the proposal loss $\mathcal{L}_{p}$ introduced in MipNeRF360 [1]. At the end of the training, we switch to computing the PDF weights directly from $\mathcal{F}_{\Theta}^{f}$ to allow the SDF representation to be freely refined.
@@
For both representations, $\mathcal{F}_{\Theta}^{v}$ and $\mathcal{F}_{\Theta}^{f}$, we employ a standard $L_{1}$ loss to minimize the pixel-wise color difference between the rendered image $\hat{C}$ and the ground truth image $\bar{C}$, along with a DSSIM [31] loss on color patches. These losses are jointly denoted as $\mathcal{L}_{\text{rgb}}$. Similar to StreetSurf [12], we model the sky color using an auxiliary MLP conditioned on the ray direction and introduce a sky loss $\mathcal{L}_{\text{sky}}$ to enforce zero opacity for sky pixels. The segmentation mask is generated using an off-the-shelf semantic segmentation network [6]. In addition to the geometric regularization described in Section. 3.3, we apply distortion regularization $\mathcal{L}_{\text{d}}$ from MipNeRF [1] to mitigate floaters. Finally, to enhance fine structure learning, we incorporate an additional semantic head into $\mathcal{F}_{\Theta}^{v}$ and impose a cross-entropy loss for semantic supervision, denoted as $\mathcal{L}_{\text{sem}}$.
@@
$\mathcal{L}_{\text{f}}=\mathcal{L}_{\text{rgb}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{\text{sky}}\mathcal{L}_{\text{sky}}+\lambda_{\hat{N}}\mathcal{L}_{N}^{u}+\lambda_{\text{eik}}\mathcal{L}_{\text{eik}}^{u},$ (11)
where $\lambda_{d}$, $\lambda_{\text{sky}}$, $\lambda_{\text{normal}}$, $\lambda_{d}$, $\lambda_{\text{normal}}$, $\lambda_{\text{eik}}$ are scaling factors. During the early epochs, we reduce the coefficients for both normal supervision and distortion regularization to facilitate a stable initialization.
@@ 4 Experiments
We use hash encoding to encode the positions [23], and spherical harmonics to encode the viewing directions. We use 2 layers with 64 hidden units for the MLPs $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$. We trained our model on a single consumer GPU with 24GB of VRAM, using Adam optimizer with a cosine learning rate decay from $10^{-2}$ to $10^{-4}$. We use Marching Cubes [19] to generate the final mesh that represents the scene. Further implementation details can be found in the supplementary materials.
@@
$\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec.
@@
$\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec. $\rightarrowM P$rec.
@@
KITTI Pandaset Waymo nuScenes
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
StreetSurf 24.04 0.83 22.24 0.66 23.42 0.77 22.28 0.76
ViiNeuS 24.83 0.89 22.95 0.80 23.74 0.87 21.96 0.83
GOF 23.34 0.86 25.54 0.87 20.92 0.81 19.13 0.76
J-NeuS (ours) 24.11 0.88 23.1 0.80 24.90 0.88 24.10 0.86
Figure 4: Mean photometric results for each dataset. We highlight best performing methods in green and second one underlined.
@@
We evaluate our method on four public driving datasets: KITTI-360 [18], nuScenes [2], Waymo Open Dataset [29], and Pandaset [39]. For each dataset, we select four diverse scenes to capture a wide range of urban settings. In the case of Pandaset, we focus on static sequences and mask dynamic vehicles to ensure consistent evaluation.
@@
We compare our proposed solution to current state-of-the-art (SoTA) surface reconstruction methods, including the SDF-based approaches StreetSurf [12] and ViiNeuS [9], as well as the Gaussian splatting-based method GoF [43]. StreetSurf [12] models close, far, and sky regions using hierarchical space partitioning, with 3D and 4D hash grids and occupancy grids for efficient ray sampling. ViiNeuS [9] initializes the SDF field with volumetric density predictions, achieving SoTA results in autonomous driving scenario. GoF [43] is a Gaussian splatting method that achieves SoTA performance in object-centric scenes with high image overlap. It leverages 2D Gaussian splatting regularization losses [14] and introduces an enhanced mesh extraction solution tailored for 3D Gaussian splatting (3DGS). In our experiments, we initialize GoF using COLMAP-derived sparse and dense point clouds, denoted as GoF-sparse and GoF-dense, respectively.
@@
In addition to conventional photometric metrics (see Table 4), we assess the quality of the reconstructed meshes using two metrics similar to those in ViiNeuS [9].
@@
  • Precision (Prec.): the percentage of LiDAR points with a distance to the mesh below 0.150.15m.
4.1 Results
Quantitative analysis.
We report quantitative results across four datasets in Tab. 1. Our method consistently outperforms or matches SoTA approaches, achieving the top metrics in most scenes on KITTI-360 and nuScenes and delivering competitive results on Pandaset and Waymo.
[…]
(a) w/o GRS
(b) w/o Ada. Relax.
(c) J-NeuS (ours)
Figure 5: Ablations study: we deactivate some J-NeuS ’s key components. (a) without GRS directed by $\mathcal{\mu}_{d,c}$ (b) without adaptive relaxation on Eikonal constrain and normal supervision
StreetSurf [12] ViiNeuS [9] GoF [43]-dense J-NeuS (ours)
@@ 4.1 Results
Figure 6: Qualitative experiments results on four popular autonomous driving datasets. Complex scene geometries reconstructed by the mesh are color highlighted. We compare our mesh to the ones from StreetSurf [12], ViiNeuS [9] and GoF [43]
@@
To complete our evaluation, we further present qualitative geometric results in Figure 6 across different dataset with complex scenes geometry, including a mixture of large road surfaces and fine structures such as poles, traffic lights, and tree trunks. Both ViiNeuS [9] and StreetSurf [12] are able to reconstruct smooth complete surface but result in incomplete fine structure reconstruction. GoF [43] produce hight quality on fine structures but fails large road plane reconstruction (see sequence 0064 of nuScenes dataset). Our results demonstrate that our method outperforms the other baselines in accurately reconstructing both surfaces and fine structures, under conditions of limited image overlap.
Efficiency
mesh extraction time train time params. size
(infer. & post-process. min) (min) (MiB)
GOF $30$ 306030-60 100300100-300
StreetSurf $\approx~1$ $40$ $92.59$
ViiNeuS <0.5<0.5 2020 $27$
J-NeuS (ours) <0.5<0.5 $35$ $51$
Table 2: Performance was evaluated on the KITTI-360 dataset using the same consumer-grade GPU with 24GB of VRAM.
Table 2 reports the computational performance of each method. J-NeuSachieves the fastest mesh extraction time—under 30s—whereas GoF requires approximately 30min. Moreover, our approach delivers the highest-quality meshes while maintaining moderate training time and memory usage, a critical combination for large-scale autonomous driving applications.
@@ 5 Conclusion
In this work, we presented J-NeuS, a novel uncertainty estimation framework that effectively combines volumetric SDF representations for robust urban scene reconstruction under limited view overlap. By estimating both photometric and geometric uncertainty, we introduce Guided Ray Sampling to deploy each representation where it excels. To avoid over regularization we propose a novel robust kernel to adaptively relax geometric regularization for SDF. Extensive quantitative and qualitative evaluations show that J-NeuS outperforms SoTA SDF and Gaussian splatting methods, delivering more accurate reconstructions on both large planar regions and fine structures under autonomous driving sensor setting.